OpenStudy (anonymous):

Create your own third degree polynomial that when divided by x + 2 has a remainder of –4. Please Help

OpenStudy (anonymous):

\[\it can be a \left( ax ^{2}+bx+c \right)\left( x+2 \right)-3,where a \neq 0\] a,b,c are integers. you can plug any value of a,b,c except a=0

OpenStudy (anonymous):

Dude Im the dumbest kid when it comes to math, can u please narrow it down for me?

OpenStudy (anonymous):

\[\left( 2x ^{2}+3x+4 \right)\left( x+2 \right)-3=2x ^{3}+3x ^{2}+4x+4x ^{2}+6x+8-3\] \[=2x ^{3}+7x ^{2}+10x+5\]

OpenStudy (anonymous):

sure that it has remainder of -4?

OpenStudy (anonymous):

i got a remainder of -3 for that polinimial

OpenStudy (anonymous):

@julian25 yeah its the question they asked me on my online class bro!

OpenStudy (anonymous):

ok man, thanks!!

OpenStudy (anonymous):

here remainder is -3 ,for -4 replace-3 by -4

OpenStudy (anonymous):

Okay, so what exactly do I put on my paper guys? :$

OpenStudy (anonymous):

maybe the problem could be easy when the problem says divide the polinomial by (x-2) and get a value of -4 we know by the remainder theorem that f(-2) =-4

OpenStudy (anonymous):

it could be any polinomial of third degree for example f(x) =ax^3

OpenStudy (anonymous):

replacing -4 = a(-2)^3 -4 =-8a a=1/2 so a possible polinomial is f(x) = (1/2)x^3

OpenStudy (anonymous):

ok so thats what i put on my paper?

OpenStudy (anonymous):

jajajjaja lol i dont understodd right

OpenStudy (anonymous):

u*

OpenStudy (anonymous):

one posible polinomial is (1/2)x^3 there is another answers but that polinomial is right